OpenStudy (anonymous):

Ian solved the system of equations by making a substitution for y in the second equation. What equation did he get as a result of that substitution? y = –3x + 1 5x – 2y = 31

OpenStudy (anonymous):

5x + 6x + 1 = 31 5x + 6x – 2 = 31 –15x + 5x = 31 –15x – x = 31